OpenStudy (miihm):

Multiply these binomials. (4x + 3)(x + 1) = How do I solve this? Can someone put it into simple steps (very simple steps)? Please and thank you!

OpenStudy (hunus):

\[\Large (a+b)(c+d)\] When you multiply two binomials, you use the distributive property twice \[\Large (a+b)(c+d)=a(c+d)+b(c+d)=ac+ad+bc+bd\]

OpenStudy (hunus):

So, for example, if you were to multiply (2x+3) by (x+7) you would get \[(2x+3)(x+7)=2x(x+7)+3(x+7)=2x^2+14x+3x+21=2x^2+15x+21\]

OpenStudy (skullpatrol):

17x

OpenStudy (miihm):

Thank you. :)

OpenStudy (skullpatrol):

that is not the answer but an error in $$2x^2+15x+21$$it should be $$2x^2+17x+21$$

OpenStudy (xanthe):

lol

OpenStudy (xanthe):